### v2

# GenTorrent: Scaling Large Language Model Serving with An Overley Network

Abstract
While significant progress has been made in research and development on open-source and cost-efficient large-language models (LLMs), serving scalability remains a critical challenge, particularly for small organizations and individuals seeking to deploy and test their LLM innovations. Inspired by peer-to-peer networks that leverage decentralized overlay nodes to increase throughput and availability, we propose GenTorrent, an LLM serving overlay that harnesses computing resources from decentralized contributors. We identify four key research problems inherent to enabling such a decentralized infrastructure: 1) overlay network organization; 2) LLM communication privacy; 3) overlay forwarding for resource efficiency; and 4) verification of serving quality. This work presents the first systematic study of these fundamental problems in the context of decentralized LLM serving. Evaluation results from a prototype implemented on a set of decentralized nodes demonstrate that GenTorrent achieves a latency reduction of over 50% compared to the baseline design without overlay forwarding. Furthermore, the security features introduce minimal overhead to serving latency and throughput. We believe this work pioneers a new direction for democratizing and scaling future AI serving capabilities.
1 Introduction
The rapid development and proliferation of open-source, cost-efficient large-language models (LLMs)–such as Google’s Gemma [2], DeepSeek-R1 [12], and Meta’s LLaMA series [43]–have significantly lowered the barrier for deploying AI systems. This progress makes it feasible to run advanced language models on personal workstations and local servers with consumer GPUs. Despite this accessibility, serving scalability remains a critical challenge for small organizations and individuals to deploy their LLM innovations. Individual researchers, university laboratories, and start-up companies often find it prohibitively expensive and technically challenging to deploy their models at scale, particularly when popularity surges quickly. For instance, some promising models initially provided excellent performance at low cost, but later suffered severe performance degradation when user demand suddenly increased [1]. Such challenges highlight the limitations of traditional cloud-based infrastructures, which are not always affordable or accessible for smaller organizations and individuals, and also pose significant privacy risks due to centralized user data management.
We propose GenTorrent, a decentralized overlay infrastructure designed for large-scale LLM deployment. Inspired by PlanetLab [11], a well-known peer-to-peer (P2P) network allowing researchers to deploy and test Internet applications across distributed nodes, GenTorrent is designed to harness computing resources from a diverse range of contributors, including individuals, academic institutions, and small organizations. Participants in GenTorrent volunteer their computing power and network resources, and in return gain the ability to deploy their own LLMs and AI-driven services on this shared, decentralized network.
GenTorrent provides the following main advantages:
  1. 1.
    Scalability, resiliency, and cost-efficiency. By distributing computational tasks across numerous decentralized nodes, GenTorrent can dynamically adapt to rapid fluctuations in demand, eliminating the performance bottlenecks commonly associated with centralized cloud deployments that depend on rigid, pre-defined service agreements. Idle computing resources (e.g., personal GPUs, university lab machines) can be utilized instead of expensive dedicated cloud instances.
  2. 2.
    Democratization of AI innovation development. GenTorrent significantly lowers the barrier to independent researchers, developers, and academic institutions being able to deploy and experiment with their LLM innovations at scale. Hence, it encourages innovation from a wider and more diverse contributor base.
  3. 3.
    Enhance user privacy. Conventional cloud deployments often centralize user prompts and associated private information, creating a single point of trust and potential vulnerability. GenTorrent addresses this by incorporating anonymous overlay networks, preventing model-serving nodes from easily linking user prompts to individual identities, thereby enhancing user privacy.
  4. 4.
    GenTorrent naturally inherits benefits that are common to decentralized overlay systems, such as service redundancy, geographic distribution of resources, and load balancing among heterogeneous servers.
GenTorrent thus represents a novel approach to deploying LLMs and other advanced AI services, democratizing large-scale AI capabilities and enhancing user privacy.
In designing and implementing GenTorrent, this work addresses four core research challenges that are critical for any decentralized LLM serving infrastructure:
  1. 1.
    Overlay network organization. We define the basic roles of overlay nodes in GenTorrent and detail the organization of the overlay network.
  2. 2.
    Anonymous communication. Decentralized model serving introduces privacy risks, as malicious overlay nodes might link user prompts to their identities. To mitigate this threat, we propose an anonymous communication service that protects user privacy by preventing the association of prompts with individual identities, even under overlay node churn.
  3. 3.
    Overlay forwarding for serving efficiency. GenTorrent enables model-serving nodes to forward user prompts across the network, allowing load balancing and key-value cache reuse. We design a decentralized forwarding mechanism that improves resource efficiency and achieves significant latency reduction while ensuring balanced workloads among model-serving nodes.
  4. 4.
    Decentralized verification of model serving. To maintain the quality and correctness of LLM serving, it is crucial to detect instances where model-serving nodes do not run their models as specified (e.g., deploying a smaller model). We develop decentralized verification to identify and mitigate such non-compliant behaviors.
To our knowledge, this is the first work to systematically study these core challenges for achieving LLM serving scalability via a decentralized overlay network independent of centralized cloud control. While further research and deployment of GenTorrent will likely reveal additional research challenges beyond the four discussed, we believe that these four issues are fundamental and critical to any decentralized LLM-serving infrastructure. We anticipate that this work will provide a foundation for future innovations and inspire continued research and development in decentralized network and system infrastructures for scalable AI serving.
We implemented GenTorrent’s core functions as a prototype consisting of decentralized nodes and tested the prototype in a public cloud without central control. This approach provided a controllable testbed, necessary as we do not have direct access to a large set of physical decentralized servers across the Internet. We let each node add synthetic latency to every packet for the wide-area Internet conditions. Upon acceptance of this paper, we will make the prototype implementation and data sets open-sourced. We plan to deploy and maintain GenTorrent at scale in the future.
This work focuses exclusively on the LLM serving problem in a decentralized overlay network and assumes model training is completed by model developers. We do not consider decentralized LLM training [48] or collaborative serving an LLM instance [5, 50, 33]. This work is also different from distributed LLM serving in a cloud with centralized scheduling [49, 41, 50, 33, 29].
The rest of this paper is organized as follows. Section 2 presents the system models and design objectives of GenTorrent. Section 3 details the design of GenTorrent. We provide the security analysis and evaluation in Section 4. The performance evaluation results are presented in Section 5. We discuss the related work in Section 6. We conclude this work and discuss possible future work in Section 7.
2 System Models and Objectives
Figure 1: Overview of GenTorrent.
This section presents the network and security models used by GenTorrent, whose overview is illustrated in Fig. 1. We also outline the design objectives of GenTorrent.
2.1 Network Model
GenTorrent operates as an overlay network composed of end hosts, each plays as one or more of the following roles:
  1. 1.
    User node represents end-users (clients) seeking access to LLM services while preserving privacy. To facilitate an anonymous overlay, each user node also agrees to forward packets for other users.
  2. 2.
    Model node executes LLM inference tasks to serve user requests. These nodes typically run on servers with GPU resources, contributed by participating organizations or individuals. Consistent with our system scope, we assume each model node serves one or more complete LLM instances. Collaborative inference across multiple nodes for a single request [5, 50] is not considered.
  3. 3.
    Verification node. To maintain correct LLM serving and defend against security attacks, GenTorrent includes a committee of decentralized verification nodes, far fewer than users and model nodes. Each verification node runs on a server owned by an organization or individual with certain credentials, such as a well-recognized university research team. They are to ensure that each model node behaves legitimately and manage user and model node information, such as handling node registration and node reputation updates for model nodes. This committee runs a BFT consensus protocol [7] to ensure information correctness and consistency.
A single physical device may operate in two or all roles in GenTorrent. All communications between nodes in GenTorrent are via TCP, secured with TLS. We assume every node is directly reachable via its IP address. User nodes may join or leave the network at any time and hence network churn can happen. Model and verification nodes are more stable but they still can fail or be compromised by attackers.
2.2 Threat Model and Assumptions
Consistent to many existing works [22, 20, 34], this work focuses on practical low-delay security and anonymity, rather than perfect security at all costs. We assume the following attacks by different nodes in GenTorrent, most of them are fundamental considerations in any decentralized model serving because they can be executed with small costs.
  • Malicious user nodes may drop packets they are responsible for forwarding. They may try to deanonymize the sender of a request and inspect packet contents. We also assume user nodes that can collude to conduct these attacks.
  • A dishonest model node may drop user requests on purpose or reply with incorrect model results to mislead users. This includes manipulating outputs or substituting the agreed-upon model with a cheaper, lower-quality one. For example, the node claims to serve an 8B-parameter LLM but runs a 1.5B model instead to conserve resources. It will try to differentiate normal requests from users and verification requests to treat them differently. Even though a model node does not perform active attacks, it might attempt to link the received prompts and their user nodes. It might falsely claim resource exhaustion to refuse service.
  • A compromised verification node could provide wrong information to users or false verification results for model nodes. We assume an adversary, potentially leveraging Sybil attacks to control multiple verification nodes. However, the adversary cannot control more than compromise more than $f$ out of $N=3f+1$ total verification nodes following the Byzantine fault tolerance (BFT) assumptions. This threshold is based on the assumption that verification nodes are run by decentralized entities, making the cost of compromising a sufficient fraction extremely high.
2.3 Design Objectives
GenTorrent is designed to achieve the following objectives:
Performance: The key performance goals include providing low-latency LLM services with high scalability. In addition, GenTorrent aims to achieve load balancing, preventing individual nodes from being overwhelmed (i.e., receiving more requests than their capacity, leading to high completion times). Inspired by efficiency gains in centralized LLM serving [49, 41, 25], GenTorrent also targets key-value (KV) cache reuse across the overlay. This introduces a novel challenge specific to the decentralized context: efficient decentralized KV cache matching.
Security, privacy, and availability: GenTorrent is designed to ensure robust service availability despite network churn, node failures, and the adversarial behaviors identified in the threat model. Although malicious users may intentionally drop packets sent by others, clients can still achieve a high probability of successful request delivery due to path redundancy. Service integrity and security are maintained by the verification committee, which detects model nodes providing incorrect or low-quality responses and maintains accurate, consistent node reputations. GenTorrent aims to preserve user anonymity and protect the confidentiality of prompt content and model responses. It is designed to prevent model nodes or other users from linking user identities (e.g., IP addresses) to their requests.
3 System Design of GenTorrent
This section details the design of GenTorrent.
3.1 Overview
The overview of GenTorrent is illustrated in Fig. 1. The design of GenTorrent includes a number of protocols that allow nodes with the same or different roles to communicate with each other. We outline the protocols associated with each node type below.
The user nodes form a dynamic overlay network. Upon joining GenTorrent, a user node registers its IP address and public key with the verification nodes, whose IP addresses and public keys are public information. It then downloads a list of active user and model nodes from a verification node, and the list is signed by more than 2/3 of the verification nodes [8]. For scalability in large deployments, verification nodes may divide the whole system into multiple regions and create a list of users and model nodes for each region. The requirement is that the number of users in the region should be sufficiently large to hide the requester’s identity, for example, $>1000$ users, and the number of model nodes should be sufficiently large to achieve load balancing. GenTorrent allows a user to use other users as relays for anonymous communication: each user request will be forwarded along an overlay path to hide the address of the original sending user, and the LLM reply will also be forwarded along the anonymous path back to the requesting user.
Each model node runs an instance of an LLM, which is deployed as a portable isolated environment, such as a container [32], provided by the LLM developer. One or more LLMs are deployed in the network, and each user request specifies which LLM it is requesting. A user request can be responded to by any model node that serves the same LLM as its request target. Model nodes use overlay forwarding to achieve load balancing and KV cache reuse among them.
The committee of verification nodes uses a recent BFT consensus protocol [7] to ensure the information maintained by them is consistent, assuming $<1/3$ of the nodes are malicious. The committee maintains a list of user nodes registered to GenTorrent, including their public keys, and a list of model nodes in GenTorrent, including their public keys and reputation scores. The reputation score of a model node quantifies the quality of its LLM responses and will be defined in Section 3.4. Verification nodes periodically request the model nodes through the anonymous overlay and update their scores based on the responses.
3.2 Anonymous Overlay of Users
The anonymous overlay of users in GenTorrent allows each user node to request the LLM service without revealing its IP address. We summarize the requirements of the anonymous overlay as follows:
  1. 1.
    GenTorrent preserves user node anonymity, which requires that other nodes cannot link a prompt with the requesting user IP. However, anonymity of the model node is unnecessary: Users can link the IPs of the model node to the responses.
  2. 2.
    The confidentiality of prompts and responses should be preserved. The contents can only be visible to the requester and model node.
  3. 3.
    The anonymous overlay should not cause high overhead on user nodes. In particular, we do not want to use public-key cryptography for prompts and responses, which introduces a large computing overhead at relay nodes.
  4. 4.
    A single anonymous path is susceptible to relay failures under network churn and traffic analysis attacks. GenTorrent uses multiple paths to provide reliable message delivery.
We analyze typical overlay anonymous communication protocols [13, 22, 20, 34], including overlay Onion routing and sliced routing [22, 34], in the Appendix and find that none of them fit the above requirements.
We propose a novel approach that achieves the advantages of both Onion and sliced routing and fits the requirements of GenTorrent exactly. In GenTorrent, each user uses Onion routing to establish $n$ proxies. In this process, path failures and redundancy do not cause high resource waste because the establishment message is very short. After proxies are established, the user and model nodes rely on Garlic for prompt and response messages.
Using S-IDA. We briefly present an existing protocol Secure Information Dispersal Algorithm (S-IDA) [23] that is essential to achieve message confidentiality, anonymous routing, and failure resilience in GenTorrent. An Information Dispersal Algorithm (IDA) splits a message $M$ into multiple fragments, each of which is in ciphertext. If the receiver gets enough fragments, it can recover the plaintext message. Two most well-known IDAs are Shamir’s secret sharing (SSS) [38] and Rabin’s IDA [35]. Both of them requires $k$ out of $n$ fragments to recover the original message $M$. However, each has its weakness. SSS ensures any $k-1$ fragments provide no information about $M$, but each fragment has the same length as $M$, results in $n\times$ bandwidth cost. In Rabin’s IDA, each fragment is around $l/k$ long where $l$ is the size of $M$ but Rabin’s IDA does not guarantee that $<k$ fragments provide no information about $M$. Secure IDA (S-IDA) [23, 34] improves the security level of IDA by combining it with symmetric encryption, such as Advanced Encryption Standard (AES). The sender applies the following steps.
1. Encrypts $M$ by an AES key $K$, getting $\{M\}_{K}$.
2. Splits $\{M\}_{K}$ into $n$ fragments, $M_{1},M_{2},...,M_{n}$ by a $k$-threshold Rabin’s IDA.
3. Splits $K$ into $n$ fragments, $K_{1},K_{2},...,K_{n}$ by a $k$-threshold SSS.
4. Constructs $n$ messages called cloves, $C_{1},C_{2},...,C_{n}$, where $C_{i}$ includes two fields $M_{i}$ and $K_{i}$.
5. Sends the cloves to the receiver using $n$ different paths.
By receiving at least $k$ cloves, the receiver:
1. Recovers $\{M\}_{K}$ using Rabin’s IDA and the key $K$ by SSS.
2. Decrypts $\{M\}_{K}$ and gets $M$.
In this way, only when a node collects $k$ cloves, it can recover the key and decrypted $\{M\}_{K}$. Without the complete key, it can only see the ciphertext related to $\{M\}_{K}$. We call the above process $(n,k)$ S-IDA.
Figure 2: Process of proxy establishment.
Figure 3: Process of querying the model node.
Figure 4: Process of sending the reply message from the model node to the user.
We present the detailed anonymous communication protocol in GenTorrent.
Step 1. Preparation. When a user node $u$ starts running the GenTorrent client program, it joins the overlay network. The verification nodes’ IP and public keys are available to the public. The user $u$ contacts an arbitrary verification node to download a list of overlay users, called the user list, and a list of model nodes, called the mode node list. The user list is a consensus among all verification nodes and signed by more than 2/3 verification nodes. Each entry in the user list includes the public key and IP address of an overlay user. The users can be further grouped into regions, each of which has a separate list including the users around a certain network location. Users in the same region will be used to form anonymous paths. Grouping users into regions will significantly reduce the latency of anonymous paths, and a region should also include a sufficient number of users (e.g., over 1000) to ensure anonymity. In addition, GenTorrent must determine two parameters $n$ and $k$ ($n>k$) for $(n,k)$ S-IDA. It allows up to $n-k$ failed paths among $n$ overlay paths.
2. Proxy and path establishment. Each user node, say $u$, needs to find $N$ proxies ($N\geq n$). To find each proxy, the user $u$ selects $l$ other users from the user list, where $l$ is the path length. Conventional wisdom indicates that anonymous paths with three hops in Tor (not including the destination) can achieve an appropriate balance between security and latency [3]. Hence GenTorrent uses $l=3$. The user $u$ then builds an Onion path using the public keys of these 3 selected users and uses the hash value of both $u$ and the last user on the path to be the path session ID $I$, as illustrated in Fig. 2. The last user $p$ on the path becomes a proxy of $u$, and every node on the path stores the predecessor and successor together with the path session ID $I$.
The above process might fail due to user dynamics but $u$ can easily try other paths and the resource loss is small, because the proxy establishment messages are short. By repeating the above process multiple times, each user $u$ can establish $N$ proxies using Onion paths. In later LLM prompts and responses, no public-key cryptographic operations are needed on the paths, because the relay users will use the stored path ID, predecessors, and successors for forwarding. In addition, different paths of $u$ has different IDs, hence colluding users on different paths cannot link the paths to the user.
3. Prompt messages. When a user node $u$ plans to send a prompt or query message $Q$ to GenTorrent, it first selects a model node from the model node list and writes its IP as the destination. As illustrated in Fig. 3, $u$ prepares $n$ slices (called cloves) of $Q$ using $(n,k)$ S-IDA and sends them to $n$ proxies respectively using the established paths. The query message $Q$ includes only the prompt and model node IP without any information about $u$. The application-layer header of each clove includes the path ID $I$ for the corresponding proxy but does not include the proxy IP. In this way, the user and proxy are still kept anonymous to the relay users. Also, the prompt messages are kept confidential. When each proxy receives the clove, it directly sends the clove to the destination model node, which is not anonymous. The prompt message also includes the IP addresses of at least $n$ proxies of $u$, which are revealed to the model nodes when it receives $\geq k$ cloves.
4. Response messages. As shown in Fig. 4, each response message is made into $n$ cloves using $(n,k)$ S-IDA. Then the model node sends $n$ cloves to $n$ proxies of $u$ respectively. Each proxy then forwards the received clove to $u$ using the pre-established anonymous path. When $u$ receives more than $k$ cloves, it can decrypt the reply message using S-IDA. The content of the response is confidential unless an attacker controls users that appear on $k$ or more of these $n$ paths. Even in that case, the attacker needs to use brute force to decode S-IDA because these $n$ paths have different IDs.
3.3 Overlay Forwarding among Model Nodes
All model nodes serving the same LLM form a logical group. Upon receiving a request from a user, a model node can serve this request or forward the request to other nodes in the same group. If the number of models in a group exceeds a size (50 in our implementation), the group can further be divided into multiple groups, potentially based on geographic regions.
The forwarding logic, illustrated in Fig. 5, is designed to consider both KV cache reusing and load balancing objectives. The detailed design includes the following components.
Hash-Radix tree. Recent studies [49, 41, 9] show that sharing and reusing KV cache among different GPUs can significantly reduce the average latency for LLM requests. However, prior work such as SGLang [49] and Preble [41] assumes a cloud environment that has a centralized scheduler or request router. This central entity receives all requests and distributes them to the appropriate GPUs. GenTorrent lacks such a central coordinator. The centralized scheduler [49, 41] uses a radix tree [30] for searching existing KV caches that can be reused for the current prompt, based on matching prefixes. Each tree node in a radix tree represents a prefix.
In a decentralized system like GenTorrent, naively requiring every model node to maintain a full copy of the radix tree built from the KV cache information of all model nodes in the group introduces significant scalability problems: 1) High memory overhead: each node needs to spend substantial memory space to store the aggregated radix tree, including the metadata for the KV cache from all other nodes; and 2) high communication overhead: each node needs to broadcast updates about the prefixes of its local KV cache to all other nodes in the group. To resolve these problems, we propose a new distributed data structure called a Hash-Radix tree (HR-tree), detailed below.
Figure 5: The logic of overlay forwarding
Figure 6: Pre-processing for an HR-tree
Figure 7: The structure of a Hash-Radix tree
The design of the HR-tree is motivated by cuckoo filters [16], which replace hash table entries with fingerprints to reduce memory cost. An HR-tree represents the aggregated KV cache state from all overlay model nodes in the group. A request prompt can be efficiently searched in the HR-tree. If the prompt matches a sufficiently long prefix path in the HR-tree, the HR-tree returns the model node that holds the KV cache corresponding to the matched prefix. Otherwise, the HR-tree reports no match.
The HR-tree construction involves a pre-processing step, shown in Fig. 6, where each prompt is divided into variable-length chunks. Each chunk is converted into a hash value computed by a universal hash function H. The example uses 8-bit hash values such as 0x0A and 0x8B. The length of each chunk $i$ in the number of tokens, $l_{i}$, is stored in an array $L$ that is computed by a deterministic algorithm called Sentry, detailed in the Appendix. The HR-tree is then built based on the sequence of hash values, as shown in Fig. 7. Each tree node stores the hash value of a chunk. Two sibling nodes indicate that their corresponding prompts share the same prefix up to the parent node but differ thereafter.
When each model node receives a user request that includes a prompt, it searches the prompt in the HR-tree. The search algorithm first applies pre-processing to divide the query prompt into chunks and compute their hashes. The algorithm then traverses a tree structure, starting from the root node. For each chunk of the querying prompt, the algorithm attempts to find a child node with the same hash value as the chunk. When there is no child node matching the hash value of the next chunk, the search stops. The HR-tree returns the search depth $d$. If $d$ exceeds a threshold, we consider it a match. Each tree node also includes a set of pointers, each referencing an entry of a separate table. The entry represents a model node that holds the corresponding KV cache of the prefix. Each entry in the table includes the node’s IP, current load (active requests), latency, and other information. HR-tree is distinct from the block hashing in vLLM [44, 24] for memory management and paging.
Similar to Bloom and cuckoo filters [4, 16], an HR-tree might also introduce false positives: a prompt that does not match a prefix can be judged as a match by the HR-tree. For a hash value with 8 bits, the probability of a hash collision is 1/256. However, a false positive requires matching $d$ tree nodes, so the false positive rate is as small as $1/256^{d}$.
State synchronization. For each model node in a group, it periodically broadcasts the hash values of the prefixes of its local KV cache, represented as a subtree of the HR-tree, as well as other local information such as active requests that are being processed or queued, and KV cache usage. Each node aggregates the subtrees and generates an HR-tree. Using compact hashes significantly reduces the synchronization cost in this step compared to radix trees that use the entire prefixes.
Load balancing is another design consideration in addition to KV cache reusing. The main objectives of load balancing are to ensure that no model node receives more requests than its capacity, and every model node can serve a certain number of requests to establish a reputation score. Following the approach in SGLang [49], the current implementation of GenTorrent uses the number of active requests queued at a model node as the load metric. To account for the inherent hardware heterogeneity in GenTorrent, we use the relative number of requests as the metric, which is defined as the number of requests divided by the score representing a model node’s hardware capacity, ranging from 1 to 10. Moreover, GenTorrent is compatible with other metrics, such as the KV cache usage levels [41].
Forwarding logic. The overlay forwarding logic, executed by every model node upon receiving a user request, determines how to forward the request, as illustrated in Fig. 5. The model node first searches the request’s prompt in its HR-tree. This search yields either a match or a miss.
If the search result is a miss, the node prioritizes load balancing. The algorithm will search its model node table, which contains the information of all model nodes in the group, collected during state synchronization. The algorithm finds the model node with the relatively fewest active requests. The user request is then forwarded to the model node.
If the search result is a match, the HR-tree identifies a list of model nodes that hold reusable KV caches. The next step is to filter this candidate list, removing any node whose current relative load is above a predefined load threshold. If there is one or more model nodes left, we pick the one with the fewest active requests and forward the current request to it. Otherwise, the logic reverts to the miss case.
Session affinity for consecutive prompts. Every model node includes its IP address in the response message. For subsequent follow-up prompts from the same user session, the user attempts to direct these prompts specifically to that designated model node, routing them via the anonymous overlay. This policy prioritizes KV cache reuse for consecutive prompts within a session.
3.4 Model verification
The committee of verification nodes is responsible for ensuring that the model nodes correctly run the models they promise to serve and for maintaining the reputation scores of the model nodes. The verification nodes run a BFT-like consensus protocol [7] to ensure data consistency and authenticity in the committee. The fundamental verification check is to compare a response from a target model node for a specific prompt against a reference result generated locally using the same LLM. However, verifying every user response is impractical due to scalability and potential user privacy issues.
GenTorrent introduces a sampling-based verification strategy. Instead of monitoring all responses, verification nodes periodically send challenge prompts to model nodes and verify their responses. These challenge prompts are routed by the anonymous overlay; hence, model nodes cannot distinguish verification probes from regular user requests. Verification nodes then assign each model node a reputation score (defined later) based on the quality of their response. This allows the system to evaluate the model node’s behavior probabilistically without compromising scalability or privacy.
Figure 8: Verification for the model node challenge response.
Each verification node locally deploys the same copy of the LLM as the model nodes. When receiving a challenge response from the model node, verification nodes evaluate it against their local reference result. The underlying principle is that, when using the same model and prompt, the conditional probability distribution over the next token is well-defined [15]. Suppose a verification node evaluates the correctness of a model node’s response $r$ containing $n$ tokens, $r=(t_{1},t_{2},\ldots,t_{n})$, to a challenge prompt. The verification node assesses the output token-by-token, as illustrated in Fig. 8. For each token $t_{k}$ at position $k$: 1) The verification node takes the prefix of the received response up to token $t_{k}$ and the original prompt. It uses its local LLM, conditioned on this input, to compute the probability distribution over all possible tokens at position $k+1$. 2) It then looks up the probability assigned by its local model to the actual token $t_{k+1}$ from the target model node. This probability is recorded. This process repeats for all tokens, generating a list of probabilities.
The verification node calculates a score for each challenge response by computing its perplexity (PPL) based on the sequence of probabilities $PPL=\exp\Bigl{(}-\frac{1}{n}\sum_{i=1}^{n}\log p(t_{i}\mid t_{<i})\Bigr{)}$. The normalized perplexity $1/PPL$ is used for a single challenge. By evaluating multiple challenges, the verification node obtains an average score $C(T)$ during time epoch $T$.
The reputations score of a model node at time epoch $T$, $R(T)$, is updated by a moving average:
$R(T)=\alpha\cdot R(T-1)+\beta\cdot C(T)\vspace{-1.5ex}$
where $R(T-1)$ is the score from the last epoch, $C(T)$ is the average score from the challenges in $T$, and $\alpha$ and $\beta$ are the weights for the moving average. In our implementation, we set $\alpha=0.4$ and $\beta=0.6$.
The punishment to the reputation for a low score should be much stronger than the reward for a high score. This ensures the model node has to respond correctly consistently to earn a high reputation. We use a sliding window approach to apply punishment to low values of $C(T)$. The verification nodes maintain a sliding window of each model including the past $W$ values of $C(T)$–the window size $W$ is set to 5 in our implementation. If $C(T)$ is smaller than a threshold $\tau$, it is considered an abnormal value. Let the count of abnormal values in the past $W$ be $c$. If $c/w$ is above a predefined threshold $\gamma$ (such as 1/5), the punishment applies as follows:
$R(T)=\alpha\cdot R(T-1)+\frac{W+1}{W+c/\gamma+2}\cdot C(T)\vspace{-1.5ex}$
Therefore, this punishment makes the reputation decrease excessively when detecting abnormal values. If a model node’s reputation drops below a critical level (like 0.4), it will be marked as untrusted.
Verification is performed epoch-by-epoch. In each epoch, the committee collaboratively assesses the correctness of model nodes and updates their reputation accordingly. GenTorrent relies on a consensus protocol Tendermint [6]. Tendermint is an open-source protocol to provide BFT consensus that ensures safety and liveness under partial synchrony. The system assumes that out of $n$ nodes, a maximum of $f$ can be Byzantine faulty, where $n\leq 3f+1$. Each update message should be signed by at least $12n/3+$ 1 nodes before commitment. In each epoch, a verification node is selected as the leader to send challenge requests to model nodes. To reduce overwhelming model nodes, only the leader of each epoch communicates with the model nodes, collects their responses, and broadcasts the results to the committee. Each verification node then locally verifies the results independently and votes on the updates.
The leader $L_{i}$ of the epoch $e_{i}$ is selected pseudo-randomly and verifiably towards the end of the previous epoch. Specifically, we use a Verifiable Random Function to select $L_{i}$ based on the final commit hash of epoch $e_{i-1}$. At the end of epoch $e_{i-1}$, the committee also agrees on the set of model nodes to be verified in epoch $e_{i}$, $M_{i}$, and the corresponding challenge prompts for each of them. This is prepared by the leader $L_{i-1}$ in this epoch. The prompt is a unique, random natural text question, indistinguishable from normal user prompts. No two model nodes should be asked the same prompt to prevent collusion or replay attacks. This pre-determined model node and prompt list is to prevent the leader $L_{i}$ from behaving maliciously in epoch $e_{i}$ by selectively ignoring certain model nodes or assigning inconsistent prompts.
In epoch $e_{i}$, the leader $L_{i}$ initiates the verification process by sending the pre-determined challenge prompts to each model node in $M_{i}$. Each model node receiving a prompt generates a response output, which is a sequence of output tokens. The model node cannot tell if this is a prompt from clients or verification nodes, because all requests are delivered through the anonymous overlay. Once responses are received, $L_{i}$ computes a credibility score for each model node and generates a reputation score accordingly using the above method. $L_{i}$ then broadcasts the signed challenge responses list and reputation scores to the committee for validation.
If the response and digital signature do not match or the model node does not respond, the leader can include an “invalid response from model node $x$” message in the broadcast message, but does not reduce $x$’s reputation. It is to prevent the leader from attacking a model node by dropping or altering the response. Only when more than $1/3$ of the committee reported invalid responses for the same model node, implying that the model node misbehaves on purpose, $x$’s reputation will be reduced.
Upon receiving the information from the leader, each verification node performs an integrity check locally. It verifies whether the list of model nodes $M_{i}$ and associated prompts match the pre-determined prompt list, and checks the signatures and timestamps of each model node response. If any inconsistency is detected - e.g., different prompts or wrong signatures - verification for the current epoch is aborted. A new leader will be selected to continue verification for the next epoch. For valid challenge responses, each verification node independently computes the credibility and reputation score using its local LLM, and compares it with the score proposed by the leader $L_{i}$. The reputation update follows Tendermint’s two-phase voting protocol: Pre-Vote and Pre-Commit. If they match with negligible variance, the verification node sends a pre-vote. If the leader receives pre-votes from more than $2/3$ of the verification nodes, it proceeds to the pre-commit phase. Again, if more than $2/3$ of the nodes pre-commit, the reputation updates are finalized and broadcast to the whole system. At the end of epoch $e_{i}$, $L_{i}$ prepares the set of model nodes $M_{i+1}$ to be verified for the next epoch and the corresponding challenge prompt list.
Potential consequences for participants deploying model nodes that achieve persistently low reputation scores include suspension or revocation of their accounts in GenTorrent.
Figure 9: Anonymity vs. fraction of
malicious nodes.
Figure 10: Confidentiality vs. fraction
of malicious nodes.
Figure 11: Credit score of different
models.
4 Security Analysis and Evaluation
This section presents the analysis and evaluation on how the proposed protocols can achieve the security properties.
4.1 Anonymity
We apply a commonly used entropy-based metric [51, 34] to measure the anonymity of a system, whose definition is presented in the Appendix. It ranges from 0 and 1, and a higher value represents better anonymity.
We use simulations to compare the normalized entropy across three systems: onion routing [36, 13], garlic cast (GC) [34], and GenTorrent, in a 10,000-node network under different fractions of malicious nodes. As shown Fig. 11, GenTorrent consistently achieves the highest level of anonymity, particularly in low-to-moderate corruption scenarios. For instance, at a corruption fraction of $f=0.05$, the normalized entropy for GenTorrent reaches $0.965$, compared to $0.954$ for onion and $0.903$ for GC. While introducing path diversity enhances resilience to individual path failures, it simultaneously increases the risk of de-anonymization if adversaries are capable of cross-path collusion. However, GenTorrent uses different path IDs for different paths, hence achieving greater robustness against colluding adversaries.
4.2 Message confidentiality
The content of a prompt or response is revealed to adversaries if the colluding adversaries exist on more than $k$ paths of the user node. where $k$ is the decoding threshold for S-IDA  [37, 42, 23]. Even if adversaries control $k$ paths, exhaustive brute-force decoding over all possible combinations of forwarded messages remains computationally prohibitive.
As shown in Fig. 11, both GenTorrent and GC show degradation in confidentiality if brute-force decoding is possible – a big assumption on adversaries’ capability – with GC dropping to $0.73$ and GenTorrent to $0.88$ under high adversarial presence ($f=10\%$). If brute-force decoding is impossible, both systems maintain almost perfect confidentiality across all scenarios.
4.3 Detecting Dishonest Model Nodes
We evaluate the effectiveness and robustness of GenTorrent’s reputation mechanism using a mixture of honest and dishonest LLMs. Model nodes that do not reply will be easily detected by the verification nodes, so we focus on detecting dishonest model nodes that use low-quality models or forged responses. Note that a model node cannot differentiate the requests from normal users and those from verification nodes, because all requests are routed through the anonymous overlay.
We use Meta-Llama-3.1-8B-Instruct-Q4_0 as the ground truth (GT) model for evaluation, representing a high-quality response baseline. To simulate dishonest or degraded behavior, we introduce four alternative models with lower capabilities or altered configurations. Specifically, m1 refers to Llama-3.2-3B-Instruct-Q4_K_M, m2 refers to Llama-3.2-1B-Instruct-Q4_K_M, m3 refers to Llama-3.2-1B-Instruct-Q4_K_S, and m4 refers to Llama-3.2-3B-Instruct-Q4_K_S. These models exhibit varying levels of performance degradation due to either reduced model size or quantization settings, and serve to evaluate the sensitivity and discriminative power of the GenTorrent verification and reputation framework.
We first run 50 prompts and evaluate the responses from all five models by measuring the credit scores in the normalized perplexity. Fig. 11 shows normalized perplexity-based credibility scores. We find that statistically, the ground truth model (GT) provides higher scores than other models.
(a) Punishment threshold $\gamma=1$.
(b) Punishment threshold $\gamma=\frac{1}{3}$.
(c) Punishment threshold $\gamma=\frac{1}{5}$.
Figure 12: Reputation changes with different punishment levels.
(a) Preparation latency
(b) Decryption latency
Figure 13: Latency of preparing cloves on model nodes and decrypting cloves
on user nodes.
Figure 14: Survival probability under churn (200 nodes/min).
Fig. 12 illustrates how the reputation scores evolve over time across 35 epochs by varying three penalty sensitivity levels: $\gamma=1$, $\frac{1}{3}$, and $\frac{1}{5}$. The results show a clear separation between GT and the weaker models, validating the effectiveness of the token-level probabilistic verification strategy. Fig. 12(a) shows the reputation trajectories under Level-1, where the penalty is relatively lenient. In this setting, the dishonest models experience slower degradation and tend to stabilize around reputation scores of 0.2 - 0.4, which may allow them to remain in the trusted pool for longer periods. With a stricter penalty in $\gamma=\frac{1}{3}$ Fig. 12(b), reputation scores of dishonest models exhibit a more significant decline, often dropping below 0.2 by period 5. For the most aggressive punishment threshold $\gamma=\frac{1}{5}$ in Fig. 12(c), results in an early and sharp decline of reputation scores, with most dishonest models falling below 0.1 within the first 5 periods.
Based on the empirical results, we set the reputation threshold to detect a dishonest model as 0.4 and $\gamma=\frac{1}{5}$.
4.4 Other Security Properties
Denial-of-Service (DoS) attacks. We consider two scenarios of DoS attacks. 1) A malicious verification node refuses to perform the verification task, either acting as the leader for an epoch or as a member. If the leader is malicious and refuses to send challenge prompts to model nodes or broadcasts challenge responses to the committee, the system simply fails the current epoch and restores the verification process in the next epoch with a new leader. If some committee members refuse to verify and vote, the system can still rely on the remaining honest verification nodes to vote and reach a consensus, which is guaranteed by BFT. 2) A model node does not respond to user requests. This behavior can be detected by the verification process.
Counterfeiting. There are three main types of counterfeiting attacks by a malicious verification leader. 1) The leader attempts to degrade some model nodes by sending them modified challenge prompts different from the pre-determined prompts agreed upon by the committee. Since model node responses always include the original prompt, other verification nodes can detect this deviation by comparing the embedded prompt in the response against the expected prompt from the pre-determined prompt list. 2) The leader alters the responses from the model nodes before broadcasting them to the committee. In GenTorrent, each response from a model node includes its digital signature, which ensures the integrity and authenticity of the responses. Any tampering by the malicious leader would be easily detected. Thus, the leader can only broadcast the original responses to the committee for scoring, but cannot alter the responses. 3) The leader can intentionally skip broadcasting responses from certain model nodes or falsely claim that a model node returned an “invalid response". To prevent this, all verification nodes reach consensus in advance on the list of model nodes to be challenged for the epoch. If verification nodes detect that some expected responses are missing or marked as invalid without justification, they will initiate their own challenge prompts to those model nodes, which are distinct from the original prompts used by the leader, to prevent auditing detection by the model nodes. If more than 2/3 of the verification nodes receive valid responses from the model nodes, the system will identify the leader as malicious. In such a case, the current epoch is aborted, and the verification process resumes in the next epoch with a newly selected leader.
5 Performance Evaluation
This section presents the performance evaluation results.
5.1 Methodology
Figure 15: Latency of LLM serving w/ and w/o HR-Tree.
We deploy the testbed of GenTorrent in a public cloud. The testbed includes eight model nodes with mid-tier hardware, each carrying one NVIDIA RTX A6000 GPU (48 GB) and serving a Meta-Llama-3 8B model, eight other model nodes with high-performance hardware, each carrying one NVIDIA A100 GPU (80 GB) and serving a DeepSeek-R1 14B model. We run two machines as the verification nodes. One carries an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, and the other carries an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD. There is no centralized coordinator in the testbed. Since we want to obtain results close to the Internet environment, we let each node delay each packet for 100ms for the Internet latency. However, we observe that the overall latency will be dominated by LLM responses, hence the Internet latency does not play a critical role. We use the most up-to-date version of vLLM [24, 45], a recently developed LLM serving library, on each single model node.
Each user request is encoded into 4 cloves routed through 3-hop paths, with successful recovery requiring at least 3 cloves ($k=3$). 8-bit hash values are used for the HR-tree to optimize the balance between memory footprint and false-positive rate. The chunk-length arrays $L$ are refreshed every 10,000 requests, and HR-tree state synchronization occurs every 5 seconds. Queries are dispatched according to a Poisson distribution with varied mean inter-arrival times, accurately simulating real-world user query patterns and request bursts commonly observed in production environments.
The evaluation includes four realistic workloads:
1. ToolUse (ToolBench[18]): Comprising 210k queries, this workload features medium-length prompts (averaging 7,206 tokens after Zipf-1.1 sampling) with tool-specific instructions. Prefix sharing is moderate, with each query capped at a 100-token output for our tested models.
2. Coding (APPS[19]): Comprising 10,000 coding problems, this workload presents coding problems followed by detailed solution requests, averaging 1,802 tokens per prompt after Zipf-0.8 sampling. Prefix overlap is minimal, and generated outputs are longer, capped at 1,000 tokens per query for our tested models.
3. Long-Doc QA (LooGLE[27]): Comprising 776 long documents and 6.4k questions. LooGLE forms long prompts, each query includes a long document as prefix and then a question about the document, averaging 10,9852 tokens per prompts after Zipf-0.6 sampling. This scenario reflects realistic document-centric query patterns, with outputs capped at 100 tokens per query.
4. Mixed Workload: Emulating realistic heterogeneous workload conditions, this workload integrates the ToolUse, Coding, and Long-Doc QA workloads in a 3:6:1 ratio[41, 47], averaging 9,959 tokens per prompts.
Upon acceptance of this paper, we will make the prototype and data sets open-sourced.
5.2 Performance of Anonymous Routing
Figure 16: Ablation study of latency.
Figure 17: Cache hit rate.
Figure 18: LLM throughput.
We evaluate the performance overhead of anonymous message (clove) processing using the Toolbench dataset [18]. Clove preparation is performed on a model node equipped with a single NVIDIA A100 40GB SXM4 GPU, 30 vCPUs, 200 GiB RAM, and 0.5 TiB SSD running Llama-3.3-70B. Decryption is executed on a user node with a single Intel i7-7700 CPU @ 3.60GHz. Figure 14 shows the cumulative distribution functions (CDFs) for both preparation and decryption latency over 10,000 trials.
Clove preparation is computationally lightweight, with a mean latency of 0.273 ms and a median (P50) latency of 0.280 ms. Even at the P99 percentile, latency remains below 0.31 ms, demonstrating tight performance bounds. Clove decryption is similarly efficient, with a slightly higher mean latency of 0.302 ms and a P50 of 0.309 ms. The P99 latency is 0.334 ms, and the success rate remains 100%, confirming the reliability of the decryption process. Overall, both operations exhibit low and stable latency.
We conduct simulations to evaluate the communication survival rate under network churn in a 3,119-node network and a churn rate of 200 nodes/min – a very high rate. A survived communication is defined as the user successfully sending the request to a model node and decoding the response by receiving at least $k$ cloves. The results are shown in Fig. 14. The survival rates of GenTorrent and GC are both much higher than that of onion. After 15 minutes, the survival rate of GenTorrent is still above 80%, without finding new proxies. Hence, we suggest each user node discover new proxies every 15 minutes.
5.3 Performance of Overlay Forwarding
We evaluate three types of latency of the model node responding to user requests: average generation latency capped at 1000 tokens (Avg), 99th percentile (P99), and Time to First Token (TTFT). The LLMs used in the evaluations are Meta Llama 3 (8B) running on A6000 and DeepSeek-R1 (14B) running on A100. Each data point reported here is the average of three independent runs of the entire workload. Latency results are shown in Fig. 15 for GenTorrent and the design without HR-tree based KV cache reusing. GenTorrent consistently demonstrates significant latency improvements compared to the design without KV cache reusing, across all evaluated workloads, LLM, and GPU setups. Notably, under heavy workload (LooGLE and Mixed), the difference is more evident, indicating the effectiveness of GenTorrent’s overlay forwarding and cache-reuse strategy. The TTFT metric directly impacts perceived user responsiveness, a critical factor in practical deployments. GenTorrent provides significantly shorter latency in TTFT. For both LLMs, GenTorrent reduces TTFT at higher request rates by 40-50%. In addition, for higher request rates, GenTorrent maintains slower latency growth.
To better understand the benefits of different design components, we conduct an ablation study by incrementally enabling the HR-tree and load balancing feature using 8 A100 as 8 model nodes running Llama-3.1-8B. The results are shown in Fig. 18 with the ToolUse workload and a Zipf-1.1 distribution. We find that using the HR-tree almost reduces both the average and P99 latency by more than half, and adding load balancing further reduces latency.
We show the KV cache hit rates of the mixed workload on 8B model nodes in Fig. 18. GenTorrent consistently achieves significantly higher cache hit rates across all evaluated workloads.Improvements in cache hit rate translate directly into throughput gains as shown in Fig. 18. We set the throughput of GenTorrent under ToolUse to 1 and show the normalized throughput. GenTorrent outperforms the design without HR-tree across all workloads.
5.4 Verification throughput
Assuming each model node requires verification 50 times per day and a deployment ratio of up to 100 model nodes per VN, the required throughput is  208 verifications per verification node per hour. We evaluate the verification nodes on two different hardware platforms. On a high-end setup equipped with an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD, a node achieves a throughput of 45.04 verifications per minute. On a standard configuration featuring an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, the verification rate is 20.72 verifications per minute. Both hardware platforms are sufficient to complete the verification tasks.
6 Related Works
With the growing importance and popularity of LLM applications, LLM serving engines, software frameworks to run LLM inferences efficiently, have been studied and developed [24, 40, 49, 26, 41, 31, 14, 39, 46, 50, 33, 29, 28, 25]. One key challenge of an LLM serving engine is the management of KV cache. For example, vLLM [24] stores KV cache as non-continuous blocks, each with a fixed number of tokens. It enables sharing KV caches for different user requests with similar prompts.
Distributed LLM serving studies the management of a group of serving nodes in a cloud/cluster environment [17, 21, 49, 41, 29]. One key research problem is to distribute requests to different serving nodes [21, 49, 41]. SGLang [49] and Preble [41] use a radix tree to find reusable KV caches across different user requests and implement scheduling algorithms to achieve load balancing. DistServe [50] assigns prefill and decoding to different nodes to eliminate their interferences. These studies assume a centralized “scheduler” or “router” that analyzes all user requests based on various scheduling algorithms and sends the requests to serving nodes. In contrast, there is no such central scheduler in GenTorrent with knowledge of the entire system. GenTorrent is the first study of the three key problems in decentralized serving.
In the literature of Internet-based LLM serving, Petals [5] proposes to run an LLM (such as BLOOM-176B) on multiple consumer GPUs collaboratively. The Knowledge Delivery Network (KDN) [10] allows storing and exchanging KV caches among different serving engines. They focus on completely different design objectives from GenTorrent.
7 Conclusion
This work introduces GenTorrent, a novel decentralized infrastructure designed for scalable and privacy-preserving LLM serving. We present designs addressing four key challenges: overlay network organization, anonymous communication, overlay forwarding for efficiency, and verification of model serving quality. These issues are fundamental and inherent to any future decentralized LLM-serving system. We implement a prototype of GenTorrent and evaluation results show that GenTorrent achieves low serving latency with minimal overhead for security and trustworthy features. In the future, we plan to deploy GenTorrent at scale and encourage broader community engagement in the research and use of GenTorrent. We believe this work offers a new direction that will attract further research and eventually help democratize future AI innovations and deployment.
Appendix
A1. Limitations of existing overlay anonymous routing.
Limitations of overlay Onion routing. Onion routing [36] was proposed to achieve anonymous routing in the network layer where the forwarding nodes are network routers, such as Tor [13]. If we apply the same idea to the overlay network in GenTorrent, the main challenge is that the failure probability of an Onion path increases exponentially with the path length, especially for an overlay network of users with a high churn rate. In addition, each relay needs to perform public-key descryption on the entire message, introducing non-trivial overhead for users.
Strengths and weaknesses of slicing protocols. One idea is to use sliced messages for overlay anonymity [22, 34]. In garlic routing [34], the sender sends $n$ slices, each is about $1/k$ of the original message size long and is called a clove. If the receiver gets $k$ out of $n$ cloves, it can recover the original message. Garlic achieves 1) anonymity without encryption and decryption operations on the paths; 2) reliability under node failures; 3) the bandwidth cost does not increase significantly with the $n-k$ redundant cloves. Garlic routing has a few weaknesses if we apply it directly to GenTorrent. 1) Garlic relies on random walks to establish routing paths and the corresponding “proxies”. However, random walks add uncertainty to the success of finding enough proxies. 2) Garlic preserves the provider (model node) anonymity, which is unnecessary for GenTorrent.
Algorithm 1 HashRadix Tree Search
1:function Search(HRTree, request, $\tau_{c}$)
2:     $d\leftarrow 0$
3:     $l_{1},l_{2},...,l_{n}\leftarrow HRTree.L$
4:     $k\leftarrow 0$
5:     for $i\leftarrow 1$ to $n$ do
6:         $chunk_{i}\leftarrow hash(request[k..k+l_{i}],HRtree.mod)$
7:         $k=k+l_{i}$
8:     end for
9:     $node\leftarrow HRTree.root$
10:     for $i\leftarrow 1$ to $n$ do
11:         if $node$ has child $chunk_{i}$ then
12:              $node\leftarrow node.children[chunk_{i}]$
13:              $d\leftarrow d+1$
14:         else
15:              return $node.modelNodeList,d$
16:         end if
17:     end for
18:     if $d<\tau_{c}$ then
19:         return $\emptyset$
20:     else
21:         return $node.modelNodeList,d$
22:     end if
23:end function
A2. Pseudocode
Algorithm 1 presents the pseudocode of the HR-Tree search algorithm.
Algorithm 2 Load Balancing
1:function LoadBalancing(HRTree, request, $\tau_{c}$)
2:     $modelNodeList\leftarrow search(HRTree,request,\tau_{c})$
3:     if $modelNodeList=\emptyset$ then $\triangleright$ Cache miss
4:         Forward to model node with fewest relative requests
5:     $\triangleright$ Cache hit
6:         $candidate\leftarrow$ Model node with fewest relative requests in $modelNodeList$
7:         if $candidate.load<candidate.threshold$ then
8:              Forward to $candidate$
9:         else
10:              Forward to one with fewest relative requests in all model nodes $\triangleright$ Fallback to load balancing
11:         end if
12:     end if
13:end function
Algorithm 2 presents the pseudocode of the load-balancing algorithm with the HR-tree result.
Algorithm 3 Model Verification
function CheckCredibility(prompt, output) $\triangleright$ Tokenized
2:     $probabilities\leftarrow[]$
     $context\leftarrow prompt$
4:     for $i\leftarrow 0$ to $length(output)-1$ do
         $next\_token\leftarrow output[i]$
6:         $logprobs\leftarrow\text{GetCompletionLogprobs}(context,max\_tokens=1)$
         $token\_prob\leftarrow\text{FindMatchingProbability}(logprobs,next\_token)$
8:         if $token\_prob$ exists then
              $probabilities.append(token\_prob)$
10:         else
              $probabilities.append(\epsilon)$triangleright {Small constant}
12:         end if
         $context\leftarrow context+next\_token$
14:     end for
     return $score\leftarrow\text{NormalizedPPL}(probabilities)$
16:end function
Algorithm 3 presents the pseudocode of the model verification process.
Figure 19: Chunk length array $L$ set by Sentries.
A3. The Sentry algrotihm for the chunk length array
Determining the optimal configuration for the chunk length array $L$ presents a significant challenge. Setting the chunk sizes too small results in a HashRadix Tree (HR-Tree) with excessive depth, potentially over-differentiating prefixes. Conversely, if chunk sizes are too large, requests with substantial prefix overlap and high cache affinity might be incorrectly treated as entirely distinct. To address this, GenTorrent incorporates a Sentry module. This component is responsible for collecting incoming requests and detecting common patterns, such as system prompts, which can inform the dynamic adjustment or initial configuration of the chunking strategy defined by $L$. How to set $L$ to effciently utlize the HR-Tree? As described in Fig. 19, we define $S=s_{1},s_{2},\dots,s_{n}$ be the lengths of the distinct common system prompts detected by the Sentries, sorted in ascending order. The Sentry module uses these lengths, along with a small, fixed separator length $\delta$, to construct the chunk length array L used in Algorithm 1. The array L is defined as follows:
$l_{1}$ $=s_{1}$ (1)
$l_{2n}$ $=\delta$ (2)
$l_{2n+1}$ $=s_{n}-s_{n-1}-\delta$ (3)
This structure ensures that each distinct system prompt corresponds to a specific path segment in the HR-Tree, separated by a small $\delta$ chunk. This allows the initial layers of the HR-Tree to be efficiently utilized for routing based on common prompts, thereby enhancing cache affinity for requests that share these initial structures. Subsequent parts of the request (beyond the detected prompts) are tokenized using any remaining lengths in $L$ or a default chunking strategy.
A4. Definition of Anonymity Metric
Definition. Let $S$ be the set of all nodes in the network with $|S|=N$. An attacker assigns to each node $x\in S$ a probability $p_{x}$ of being the source (or destination) of a message. The entropy of the system is defined as:
$H(S)=-\sum_{x\in S}p_{x}\log_{2}(p_{x}).$
Perfect anonymity occurs when the attacker has no information about the system, and thus must assign a uniform probability of $1/N$ to each node. In this case, the entropy reaches its maximum:
$H_{\max}=\log_{2}(N).$
To evaluate the anonymity of the system, we normalize the entropy as a metric. The normalized anonymity is given by:
$\frac{H(S)}{H_{\max}}=\frac{-\sum_{x\in S}p_{x}\log_{2}(p_{x})}{\log_{2}(N)}.$
We use the above metric but need to know how to compute $p_{x}$. We assume a fraction $f$ of all users are collaborative malicious users. Some of them may sit on the $k$ paths between the client and the proxies. However, since they have no information on the prompt, they cannot tell if they are serving the same client. Suppose the number of all nodes on the $k$ paths is $L$. There may exist multiple chains of consecutive attackers on the paths. The attackers may guess that the predecessor of every chain is the source and the successor of every chain is the destination. Hence the probability of a guess being correct is $\frac{1}{L+1-f\cdot L}$. Let $\Gamma$ be the set of predecessors of all malicious chains. We then have,
$\Pr(x=\text{src})=\begin{cases}\frac{1}{L+1-f\cdot L}&\text{if }x\in\Gamma\\ (1-\frac{|\Gamma|}{L+1-f\cdot L})\frac{1}{(1-f)N-|\Gamma|}&\text{otherwise}\end{cases}$
The entropy-based anonymity can be computed using the above probabilities.